Opened 10 years ago

Last modified 5 years ago

#402 assigned flight processing

CAM11_03, flight day 141/2011, Madrid — at Version 11

Reported by: knpa Owned by: iopa
Priority: alpha 4 high Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by anch)

Data location: /users/rsg/arsf/arsf_data/2011/flight_data/spain_portugal/CAM11_03-2011_141_Madrid

Data arrived from ARSF via network transfer on 24/05/2011

CAM11_03 Scientific objective: We wish to measure how these forests change over gradients of topography‐controlled climate and soil properties.

CAM11_03 Priority: Alpha 4h

PI: David Coomes

Sensors:

Camera (requested)
Eagle (27/06/2011)
Hawk (27/06/2011)
LiDAR (requested)

Note: This project is actually part of EU11_03 but has a different project code as it was funded separately, and has higher priority.

Change History (11)

comment:1 Changed 10 years ago by knpa

  • Description modified (diff)

comment:2 Changed 10 years ago by iopa

Taking LiDAR and camera.

comment:3 Changed 10 years ago by iopa

Camera is ready for checking @ ~arsf/arsf_data/2011/flight_data/spain_portugal/CAM11_03-2011_141_Madrid/processing/delivery/camera-CAM11_03-100-20110615

Raw images taken during preparation and removed ;
1-28
Raw images with cloud or saturation and removed ;
155,194,276,607-615,699

comment:4 Changed 10 years ago by knpa

  • Milestone set to 2011 data processing completion
  • Owner set to iopa
  • Status changed from new to assigned

Taking Camera delivery check.

comment:5 Changed 10 years ago by knpa

Delivery check complete. All looks good apart from:

  1. Looks like the calibration certificate from last year. Is that OK?
  2. Noticed in the tiff tags that the gps time is given as 10/04/11, not 10/05/11 when the project was flown.

comment:6 Changed 10 years ago by benj

LiDAR processing currently blocked until we get a new boresight flight

comment:7 Changed 10 years ago by anch

Hyperspectral ready for delivery check

comment:8 Changed 10 years ago by knpa

Taking hyperspectral delivery check.

comment:9 Changed 10 years ago by knpa

Delivery check complete.

All looks good apart from:

  1. Logsheet top section has lots missing. Can fill this out from the scanned logsheet.
  2. Readme: aplcorr command needs altering. e141011b_post_processed_nav.bil should be e141011b_nav_post_processed.bil (script needs editing also)
  3. Blank bands at 127 and 0 (eagle)
  4. Blank band at 0 (hawk)

comment:10 Changed 10 years ago by anch

Issues with delivery resolved.

Now putting the delivery on a disk

comment:11 Changed 10 years ago by anch

  • Description modified (diff)

Hawk and Eagle data sent to PI on 27/06/2011

Note: See TracTickets for help on using tickets.